Tyler buys 3 pieces of machinery from the machine store. Because he bought 3 pieces in a bundle he was provided with a discount of $3,500 on the entire purchase. What is his cost basis in each machine?

Machine 1: $5,000

Machine 2: $10,000

Machine 3: $7,500
Answer #1
Total Cost of Machinery= 5000+10000+7500=22500
Total Discount = $3500
Cost basis for Machine-1= Machine Cost - Allocated Depreciation
={ $5000-   (3500/22500)X 5000}=4222
Cost basis for Machine-2= Machine Cost - Allocated Depreciation
={ $10000-   (3500/22500)X 10000}=$8444
Cost basis for Machine-3= Machine Cost - Allocated Depreciation
={ $7500-   (3500/22500)X 7500}=$6333